An example of the exam class

Exam, Form: A Name: Student Number: TA: Date:

Section 1. Musicology 101

Match the following artists with the listed albums.

Andrew Lloyd Weber Jesus Jones Love and Rockets The Dave Matthews Band Nirvana Soundgarden The Jesus and Mary Chain Elvis Costello Pink Floyd The Cult Ned’s Atomic Dustbin (a) Sonic Temple

(b) The Wall

(c) Spike

(d) Doubt

(e) Nevermind

(f) Soundtrack to ‘Cats’

(g) Godfodder

(h) Earth, Sun, Moon

(i) Automatic

(j) Louder than Love

(k) Under the Table and Dreaming

True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do.

All animals are created equal, but some animals are more equal than others.

I don’t know why you’d need this, but here is a block of true/false questions.

‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. Laden swallows fly faster than unladen swallows, unless they carry coconuts. This sentence is not false. ‘Monty Python and the Holy Grail’ is a very funny movie.

Section 3. Fill in the blank (5 pts each)

Here is a place where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. Mill’s theory of morality is known as

2. How much  would a  chuck, if a  would  , wood?

3. According to Kant, we should always always always follow the imperative.

4.  ’s first work was the Tractatus-  Philosophicus.

I don’t know why you’d need this, but here is a block of fill-in-the-blank questions.

5. thought that without a strong, , effective government, chaos would reign in the state of nature.

6. One main component of Nietzche’s moral philosophy is the .

Section 4. A title

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines). Here is a block of multiple choice questions.

7.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 8.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Martin Heidegger

(d) Jacques Cousteau

9.

This is a question. Which is the correct answer?

(a) (b) (c) (d) (e)

10.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 11.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four 12.

How many elves does it take to mine tin?

(a) One.

(b) Two.

(c) Three.

(d) Four

Section 5. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

13. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

14. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

15. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

16. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

17. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

18. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

19. Here is an equation:

f(x) = h(x)  (1)
g(x) = j(x)  (2)

‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

This is\some\verbatim \text $!@#$!@#%!@#
\Yes!

20. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

21. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

22. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

23. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

24. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

The first question is 25 and the last is 26. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.

25. As you saw in equation in question— ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’

26. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought.

27. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

Answer Key for Exam A

Section 1. Musicology 101

Match the following artists with the listed albums. \underline{(f)} Andrew Lloyd Weber \underline{(d)} Jesus Jones \underline{(h)} Love and Rockets \underline{(k)} The Dave Matthews Band \underline{(e)} Nirvana \underline{(j)} Soundgarden \underline{(i)} The Jesus and Mary Chain \underline{(c)} Elvis Costello \underline{(b)} Pink Floyd \underline{(a)} The Cult \underline{(g)} Ned’s Atomic Dustbin (a) Sonic Temple

(b) The Wall

(c) Spike

(d) Doubt

(e) Nevermind

(f) Soundtrack to ‘Cats’

(g) Godfodder

(h) Earth, Sun, Moon

(i) Automatic

(j) Louder than Love

(k) Under the Table and Dreaming

True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do. \overline{False} All animals are created equal, but some animals are more equal than others. I don’t know why you’d need this, but here is a block of true/false questions. \overline{True} ‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. \overline{False} Laden swallows fly faster than unladen swallows, unless they carry coconuts. \overline{True} This sentence is not false. \overline{True} ‘Monty Python and the Holy Grail’ is a very funny movie.

Section 3. Fill in the blank (5 pts each)

Here is a place where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. Mill’s theory of morality is known as Utilitarianism

2. How much wood would a woodchuck chuck, if a woodchuck would chuck, wood?

3. According to Kant, we should always always always follow the categorical imperative.

4. Wittgenstein’s first work was the Tractatus-Logico Philosophicus.

I don’t know why you’d need this, but here is a block of fill-in-the-blank questions.

5. Hobbes thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

6. One main component of Nietzche’s moral philosophy is the will to power.

Section 4. A title

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines). Here is a block of multiple choice questions.

7.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 8.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Martin Heidegger

(d) Jacques Cousteau

9.

This is a question. Which is the correct answer?

(a) (b) (c) (d) (e)

10.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 11.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four 12.

How many elves does it take to mine tin?

(a) One.

(b) Two.

(c) Three.

Section 5. Short Answer (10 pts each)

(d) Four This is an example of the shortanswer type of question, where the questions are rearranged between tests.

13. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

14. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

15. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

16. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

17. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

18. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

19. Here is an equation:

f(x) = h(x)  (1)
g(x) = j(x)  (2)

‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

This is\some\verbatim \text $!@#$!@#%!@#
\Yes!

Ask Lewis Carroll.

20. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

21. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

22. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

23. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

24. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

The first question is 25 and the last is 26. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.

25. As you saw in equation in question— ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’ Ask Lewis Carroll.

26. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought. Ask Lewis Carroll.

27. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

An example of the exam class

Exam, Form: B Name: Student Number: TA: Date:

Section 1. Musicology 101

Match the following artists with the listed albums.

The Cult Nirvana The Dave Matthews Band Andrew Lloyd Weber Jesus Jones Ned’s Atomic Dustbin Love and Rockets Pink Floyd The Jesus and Mary Chain Soundgarden Elvis Costello (a) Nevermind

(b) Spike

(c) Godfodder

(d) Sonic Temple

(e) Doubt

(f) Louder than Love

(g) Under the Table and Dreaming (h) The Wall

(i) Earth, Sun, Moon

(j) Soundtrack to ‘Cats’

(k) Automatic

True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do.

All animals are created equal, but some animals are more equal than others. This sentence is not false.

I don’t know why you’d need this, but here is a block of true/false questions.

‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. Laden swallows fly faster than unladen swallows, unless they carry coconuts. ‘Monty Python and the Holy Grail’ is a very funny movie.

Section 3. Fill in the blank (5 pts each)

Here is a place where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. How much would a chuck, if a would , wood?

2. According to Kant, we should always always always follow the imperative.

3.  ’s first work was the Tractatus-  Philosophicus.

I don’t know why you’d need this, but here is a block of fill-in-the-blank questions.

4. thought that without a strong, , effective government, chaos would reign in the state of nature.

5. One main component of Nietzche’s moral philosophy is the .

6. Mill’s theory of morality is known as

Section 4. A title

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines). Here is a block of multiple choice questions.

7.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 8.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Martin Heidegger

(d) Jacques Cousteau 9.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 10.

How many elves does it take to mine tin?

(a) One.

(b) Two.

(c) Three.

(d) Four 11.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four 12.

This is a question. Which is the correct answer?

(a)

(b)

(c)

(d)

(e)

Section 5. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

13. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

14. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

15. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

16. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

17. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

18. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

19. Here is an equation:

f(x) = h(x)  (1)
g(x) = j(x)  (2)

‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

This is\some\verbatim \text $!@#$!@#%!@#
\Yes!

20. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

21. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

22. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

23. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

24. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

25. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

The first question is 26 and the last is 27. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.

26. As you saw in equation in question— ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’

27. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought.

Answer Key for Exam B

Section 1. Musicology 101

Match the following artists with the listed albums.

\underline{(d)} The Cult

\underline{(a)} Nirvana

\underline{(g)} The Dave Matthews Band \underline{(j)} Andrew Lloyd Weber \underline{(e)} Jesus Jones \underline{(c)} Ned’s Atomic Dustbin \underline{(i)} Love and Rockets \underline{(h)} Pink Floyd \underline{(k)} The Jesus and Mary Chain \underline{(f)} Soundgarden \underline{(b)} Elvis Costello (a) Nevermind

(b) Spike

(c) Godfodder

(d) Sonic Temple

(e) Doubt

(f) Louder than Love

(g) Under the Table and Dreaming (h) The Wall

(i) Earth, Sun, Moon

(j) Soundtrack to ‘Cats’

(k) Automatic

True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do. \overline{False} All animals are created equal, but some animals are more equal than others. \overline{True} This sentence is not false. I don’t know why you’d need this, but here is a block of true/false questions. \overline{True} ‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. \overline{False} Laden swallows fly faster than unladen swallows, unless they carry coconuts. \overline{True} ‘Monty Python and the Holy Grail’ is a very funny movie.

Section 3. Fill in the blank (5 pts each)

Here is a place where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. How much wood would a woodchuck chuck, if a woodchuck would chuck, wood?

2. According to Kant, we should always always always follow the categorical imperative.

3. Wittgenstein’s first work was the Tractatus-Logico Philosophicus. I don’t know why you’d need this, but here is a block of fill-in-the-blank questions.

4. Hobbes thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

5. One main component of Nietzche’s moral philosophy is the will to power.

6. Mill’s theory of morality is known as Utilitarianism

Section 4. A title

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines). Here is a block of multiple choice questions.

7.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 8.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Martin Heidegger

(d) Jacques Cousteau 9.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 10.

How many elves does it take to mine tin?

(a) One.

(b) Two.

(c) Three.

(d) Four 11.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four

12.

This is a question. Which is the correct answer?

(a) (b) (c) (d) (e)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

13. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

14. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

15. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

16. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

17. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

18. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

19. Here is an equation:

f(x) = h(x)  (1)
g(x) = j(x)  (2)

‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

This is\some\verbatim \text $!@#$!@#%!@#
\Yes!

Ask Lewis Carroll.

20. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

21. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

22. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

23. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

24. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

25. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

The first question is 26 and the last is 27. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.

26. As you saw in equation in question— ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’ Ask Lewis Carroll.

27. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought. Ask Lewis Carroll.

An example of the exam class

Exam, Form: C Name: Student Number: TA: Date:

Section 1. Musicology 101

Match the following artists with the listed albums.

Love and Rockets Jesus Jones Ned’s Atomic Dustbin The Dave Matthews Band The Cult Pink Floyd Nirvana Elvis Costello Soundgarden Andrew Lloyd Weber The Jesus and Mary Chain (a) Sonic Temple

(b) Nevermind

(c) Under the Table and Dreaming (d) The Wall

(e) Louder than Love

(f) Doubt

(g) Godfodder

(h) Automatic

(i) Soundtrack to ‘Cats’

(j) Earth, Sun, Moon

(k) Spike

True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do.

‘Monty Python and the Holy Grail’ is a very funny movie.

I don’t know why you’d need this, but here is a block of true/false questions.

‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. Laden swallows fly faster than unladen swallows, unless they carry coconuts. All animals are created equal, but some animals are more equal than others. This sentence is not false.

Section 3. Fill in the blank (5 pts each)

Here is a place where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about. I don’t know why you’d need this, but here is a block of fill-in-the-blank questions.

1. thought that without a strong, , effective government, chaos would reign in the state of nature.

2. One main component of Nietzche’s moral philosophy is the .

3. According to Kant, we should always always always follow the imperative.

4.  ’s first work was the Tractatus-  Philosophicus.
5. How much  would a  chuck, if a  would  , wood?

6. Mill’s theory of morality is known as

Section 4. A title

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

7.

How many elves does it take to mine tin?

(a) One.

(b) Two.

(c) Three.

(d) Four Here is a block of multiple choice questions.

8.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 9.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Martin Heidegger

(d) Jacques Cousteau 10.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 11.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four 12.

This is a question. Which is the correct answer?

(a)

(b)

(c)

(d)

(e)

Section 5. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

13. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

14. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

15. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

16. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

17. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

18. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

19. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

20. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

21. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

22. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

23. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

24. Here is an equation:

f(x) = h(x)  (1)
g(x) = j(x)  (2)

‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

This is\some\verbatim \text $!@#$!@#%!@#
\Yes!

The first question is 25 and the last is 26. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.

26. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought.

27. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

Answer Key for Exam C

Section 1. Musicology 101

Match the following artists with the listed albums.

\underline{(j)} Love and Rockets \underline{(f)} Jesus Jones \underline{(g)} Ned’s Atomic Dustbin \underline{(c)} The Dave Matthews Band \underline{(a)} The Cult \underline{(d)} Pink Floyd \underline{(b)} Nirvana \underline{(k)} Elvis Costello \underline{(e)} Soundgarden \underline{(i)} Andrew Lloyd Weber \underline{(h)} The Jesus and Mary Chain (a) Sonic Temple

(b) Nevermind

(c) Under the Table and Dreaming (d) The Wall

(e) Louder than Love

(f) Doubt

(g) Godfodder

(h) Automatic

(i) Soundtrack to ‘Cats’

(j) Earth, Sun, Moon

(k) Spike

True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do. \overline{True} ‘Monty Python and the Holy Grail’ is a very funny movie. I don’t know why you’d need this, but here is a block of true/false questions. \overline{True} ‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. \overline{False} Laden swallows fly faster than unladen swallows, unless they carry coconuts. \overline{False} All animals are created equal, but some animals are more equal than others. \overline{True} This sentence is not false.

Section 3. Fill in the blank (5 pts each)

Here is a place where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about. I don’t know why you’d need this, but here is a block of fill-in-the-blank questions.

1. Hobbes thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

2. One main component of Nietzche’s moral philosophy is the will to power.

3. According to Kant, we should always always always follow the categorical imperative.

4. Wittgenstein’s first work was the Tractatus-Logico Philosophicus.

5. How much wood would a woodchuck chuck, if a woodchuck would chuck, wood?

6. Mill’s theory of morality is known as Utilitarianism

Section 4. A title

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

7.

How many elves does it take to mine tin?

(a) One.

(b) Two.

(c) Three.

(d) Four Here is a block of multiple choice questions.

8.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 9.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Martin Heidegger

(d) Jacques Cousteau 10.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 11.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four

12.

This is a question. Which is the correct answer?

(a) (b) (c) (d) (e)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

13. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

14. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

15. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

16. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

17. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

18. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

19. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.

20. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Ask Lewis Carroll.

21. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

22. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Ask Lewis Carroll.

23. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Ask Lewis Carroll.

24. Here is an equation:

f(x) = h(x)  (1)
g(x) = j(x)  (2)

‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

This is\some\verbatim \text $!@#$!@#%!@#
\Yes!

Ask Lewis Carroll.

The first question is 25 and the last is 26. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.

25. As you saw in equation in question— ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’ Ask Lewis Carroll.

26. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought. Ask Lewis Carroll.

27. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Ask Lewis Carroll.